OpenStudy (anonymous):

f you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation. 3x + y + 2z = 1 x + 2y - z = -8 3x - y - 3z = 18

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

So, in the second equation, as instructed, isolate x IE rearrange it so that x stands alone in one side of the equation.

OpenStudy (anonymous):

a. 7y - z = 25 5y - 6z = 42 b. -5x - 5z = -10 6x - z = 19 c. 7x + 3z = -10 0x - 5z = 19 d. -5y + 5z = 25 -7y = 42

OpenStudy (anonymous):

x=-8-2y+z

OpenStudy (anonymous):

\[x=-8-2y+z\]

OpenStudy (anonymous):

@Turner

terenzreignz (terenzreignz):

That's right. Now replace the x in both the first and third equations by this new expression.

OpenStudy (anonymous):

alright one sec

OpenStudy (anonymous):

\[3(-8-2y+2)+2z=1\]

OpenStudy (anonymous):

\[3(-8-2y+z)-y-3z=18\]

OpenStudy (anonymous):

the 2 should be a z in the ( )

terenzreignz (terenzreignz):

So, simplify this. Do the same with the third equation.

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

-24-6y+5z=1

OpenStudy (anonymous):

-24-5y+z=18

OpenStudy (anonymous):

is that correct

terenzreignz (terenzreignz):

Check if you're using the correct equation! 3x + y + 2z = 1 Now, do it again, replace the x.

OpenStudy (anonymous):

ok

terenzreignz (terenzreignz):

x = z - 2y - 8

OpenStudy (anonymous):

so it should look like this -24-6y+3z+y+2z=1

OpenStudy (anonymous):

-24-7y+5z=1

terenzreignz (terenzreignz):

-6y + y = 7y? recheck.

OpenStudy (anonymous):

-24-5y+5z=1

terenzreignz (terenzreignz):

Yes. But bring that -24 over to the other side.

OpenStudy (anonymous):

-5y+5z=25

terenzreignz (terenzreignz):

I guess at this point we can stop. Only one of the choices has this equation. You can verify that doing the same substitution to the third equation will add up.

OpenStudy (anonymous):

ok i have one more

OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and second equations, then the second and third equations. x + y - 3z = -8 2x + 2y + z = 12 3x + y - z = -2